# On the Localization of the Christoffel-Darboux Kernel

### v1

@@ 2 Christoffel-Darboux measures put no mass on subvarieties
For any subvariety $Y\subset X$ and any $\epsilon>0$, there exists an open neighborhood $U$ of $Y$, so that for any $k\in\mathbb{N}$, we have
@@
Note that equilibrium measure puts no mass on pluripolar subsets, when $\mu$ is Bernstein-Markov. Hence, if $\mu$ is Bernstein-Markov, Theorem 2.1 follows from the equidistribution result of [3] showing the weak convergence of $\mu_{k}^{\rm{Berg},\Delta}$ towards the equilibrium measure.
In order to prove Theorem 2.1, we need to consider the so-called partial Christoffel-Darboux kernels associated with $Y$, denoted by $B_{k}^{\epsilon,Y}(x,y)\in L^{k}_{x}\otimes(L^{k}_{y})^{*}$ for $\epsilon>0$, and defined as
@@
where $s_{i}$, $i=1,\ldots,n_{k}(\epsilon)$ is an orthonormal basis of $(H^{0}(X,L^{\otimes k}\otimes\mathcal{J}_{Y}^{\lceil\epsilon k\rceil}),{\textrm{Hilb}}_{k}(h^{L},\mu))$, $n_{k}(\epsilon)\mathrel{\mathop{\ordinarycolon}}=\dim H^{0}(X,L^{\otimes k}\otimes\mathcal{J}_{Y}^{\lceil\epsilon k\rceil})$, and $\mathcal{J}_{Y}$ is the sheaf of holomorphic sections vanishing along $Y$. The following result, out of independent interest, will be established in Section 3.
@@
For any $\epsilon,\delta>0$, there is an open neighborhood $U$ of $Y$, so that for any $x\in U$, $k\in\mathbb{N}$, we have $|B_{k}^{\epsilon,Y}(x,x)|\leq\exp(-\delta k)$.
@@
When $\mu$ is a volume form, Theorem 2.3 follows from more precise estimates of Berman [2, Lemma 4.1].
@@
By Fujita’s theorem [16], the above upper limit is actually a limit. The definition of the volume functional naturally extends to the $\mathbb{Q}$-line bundles, and the resulting functional extends continuously to the full Néron-Severi space $NS^{1}(X)$, cf. [21, Theorem 2.2.44].
@@
For the choice $\epsilon\mathrel{\mathop{\ordinarycolon}}=\epsilon^{\prime}$ and $\delta\mathrel{\mathop{\ordinarycolon}}=1$, we now consider a neighborhood $U$ of $Y$ given by Theorem 2.3, so that for any $x\in U$, $k\in\mathbb{N}$, we have
$|B_{k}^{\epsilon^{\prime},Y}(x,x)|\leq\exp(-k).$ (2.6)
By writing $B_{k}(x,x)=\big{(}B_{k}(x,x)-B_{k}^{\epsilon^{\prime},Y}(x,x)\big{)}+B_{k}^{\epsilon^{\prime},Y}(x,x)$, and using the elementary identities
@@
together with (2.5), we deduce the bound (2.1) for such $U$ directly from (2.6). ∎
3 Forbidden region for the partial Christoffel-Darboux kernels
The main goal of this section is to establish that the partial Christoffel-Darboux kernels decrease exponentially in the neighborhood of the subvariety, that is, to prove Theorem 2.3. The proof relies on the use of several psh envelopes constructions.
More specifically, we reduce first the estimate of the Christoffel-Darboux kernel associated with the $L^{2}(\mu)$-norm to a related estimate for the $L^{\infty}$-norm. For this, we rely on the non-negligible psh envelopes introduced in [18] and further studied in [15]. Then we apply the tautological maximum principle to relate this estimate with psh envelopes associated with divisors introduced in [2]. Theorem 2.3 then follows from some estimates on such envelopes.
To simplify the presentation, we claim that it suffices to establish Theorem 2.3 under the assumption that $\mathcal{J}_{Y}=\mathscr{O}_{\hat{X}}(-E)$ for some effective divisor $E$ on $X$. To see this, let $\pi\mathrel{\mathop{\ordinarycolon}}\hat{X}\to X$ denote the blow-up of $X$ along $\mathcal{J}_{Y}$, and let $\mathscr{O}_{\hat{X}}(-E)\mathrel{\mathop{\ordinarycolon}}=\pi^{*}\mathcal{J}_{Y}$. If $\hat{X}$ is not smooth, we perform further blow-ups to resolve its singularities. By an abuse of notation, we denote the resolution by $\hat{X}$. Let $\hat{\mu}$ be a measure on $\hat{X}$ such that $\pi_{*}\hat{\mu}=\mu$. Since $\mu$ is non-pluripolar and $\pi$ is a biholomorphism outside an analytic (hence pluripolar) subset, the measure $\hat{\mu}$ is uniquely defined and remains non-pluripolar. The isomorphism (2.4) is then an isometry with respect to the induced $L^{2}$-norms. In particular, the associated diagonal Christoffel-Darboux measures $\hat{\mu}_{k}^{\mathrm{Berg},\Delta}$ on $\hat{X}$ and $\mu_{k}^{\mathrm{Berg},\Delta}$ on $X$ satisfy the compatibility relation $\pi_{*}\hat{\mu}_{k}^{\mathrm{Berg},\Delta}=\mu_{k}^{\mathrm{Berg},\Delta}$. It follows that it is enough to prove Theorem 2.3 on $\hat{X}$. From now on, we assume $\mathcal{J}_{Y}=\mathscr{O}_{\hat{X}}(-E)$ for some effective divisor $E$ on $X$.
Our proof of Theorem 2.3 will be based on the following three auxiliary statements. For the first one, we use the following notation: for a continuous metric $h^{L}_{0}$ on $L$, we denote by ${\textrm{Ban}}_{k}^{\infty}(h^{L}_{0})$ the $L^{\infty}$-norm on $H^{0}(X,L^{\otimes k})$ associated with $h^{L}_{0}$. We also denote $\omega\mathrel{\mathop{\ordinarycolon}}=2\pi c_{1}(L,h^{L})$.
Lemma 3.1.

### v2

@@ 2.3 Reduction to measures supported away from subvarieties
For any Borel measure $\mu$ which does not give full mass to pluripolar subsets, any subvariety $Y\subset X$ and any $\epsilon>0$, there exists an open neighborhood $U$ of $Y$ and $k_{0}\in\mathbb{N}$, so that for any $k\geq k_{0}$, we have
@@
To formulate a general form of the continuity principle, we fix a closed subset $A\subset X$ and consider a decreasing sequence of open neighborhoods $U_{1}\Supset U_{2}\Supset\cdots$ of $A$ such that $\bigcap_{i=1}^{+\infty}U_{i}=A$. Such a sequence always exists since $A$ is closed. By Urysohn’s lemma, we consider bump functions $\rho_{i}\mathrel{\mathop{\ordinarycolon}}X\to[0,1]$, $i\in\mathbb{N}^{*}$, verifying $\rho_{i}=1$ over $X\setminus U_{i}$ and $\rho_{i}=0$ over $U_{i+1}$. We assume that for any $\epsilon>0$, there is $i\in\mathbb{N}^{*}$, so that for any $k\in\mathbb{N}$, we have
$\int_{U_{i}}\mu_{k}^{\rm{Berg},\Delta}\leq\epsilon.$ (2.21)
We denote $\mu_{i}\mathrel{\mathop{\ordinarycolon}}=\rho_{i}\cdot\mu$, and let $B_{k,i}(x,y)\in L^{k}_{x}\otimes(L^{k}_{y})^{*}$, be the Bergman kernel associated with $h^{L}$ and $\mu_{i}$. We denote by $\mu_{k,i}^{\rm{Berg}}$ the associated probability measures on $X\times X$, defined as in (1.5). The main result of this section goes as follows.
Proposition 2.4.
Lemma 2.5.
The proof of Lemma 2.5 is based on the following comparison result on the off-diagonal Bergman kernels due to Lubinsky [41, (3.3)].
Lemma 2.6.
Assume that the measures $\mu_{1}$ and $\mu_{2}$ are ordered as $\mu_{1}\leq\mu_{2}$. Then for any $x\in X$, the associated Bergman kernels are related as
@@
Observe that the characterization of the Bergman kernel analogous to (3.23) directly ensures that the right-hand side of (2.23) is positive.
Proof.
Lubinsky in [41] establishes (2.23) only in the setting of measures on the real line, but the argument relies exclusively on the reproducing property of the Bergman kernel, and so it remains valid in our generality. For the convenience of the reader, we reproduce the argument below.
Using the fact that the Bergman kernel is self-adjoint, we write
$\Big|B_{k,1}(x,y)-B_{k,2}(x,y)\Big|_{(h^{L})^{k}}^{2}=\Big(B_{k,1}(x,y)-B_{k,2}(x,y)\Big)\cdot\Big(B_{k,1}(y,x)-B_{k,2}(y,x)\Big).$ (2.24)
Now, by the reproducing property of the Bergman kernel, we have
$\int B_{k,1}(x,y)\cdot B_{k,1}(y,x)\cdot d\mu_{1}(y)=B_{k,1}(x,x),$ (2.25)
$\int B_{k,1}(x,y)\cdot B_{k,2}(y,x)\cdot d\mu_{1}(y)=B_{k,2}(x,x).$
By taking the adjoint of the second equation above, we have
$\int B_{k,2}(x,y)\cdot B_{k,1}(y,x)\cdot d\mu_{1}(y)=B_{k,2}(x,x).$ (2.26)
Finally, by using the bound $\mu_{1}\leq\mu_{2}$, we obtain
$\int|B_{k,2}(x,y)|_{(h^{L})^{k}}^{2}\cdot d\mu_{1}(y)\leq\int|B_{k,2}(x,y)|_{(h^{L})^{k}}^{2}\cdot d\mu_{2}(y),$ (2.27)
which we rewrite using the reproducing property of the Bergman kernel and the fact that the Bergman kernel is self-adjoint as
$\int B_{k,2}(x,y)\cdot B_{k,2}(y,x)\cdot d\mu_{1}(y)\leq B_{k,2}(x,x).$ (2.28)
A combination of the above estimates yield (2.23). ∎
Proof of Lemma 2.5.
By our definition of $\mu_{i}$, we have the following bound
Remark, however, that we have
$\int B_{k,i}(x,x)\cdot d\mu_{i}(x)=n_{k},$ (2.31)
and similarly for the Bergman kernel associated with $\mu$. By this and (2.21), for any $\epsilon>0$, there are $i_{0},k_{0}\in\mathbb{N}$, so that for any $i\geq i_{0}$, $k\geq k_{0}$,
Proof of Proposition 2.4.
Note, however, that by the definition of $\mu_{i}$, we have
A combination of (2.33), (2.34) and (2.35) yields Proposition 2.4. ∎
2.4 Decay of the kernel for measures supported away from a divisor
This section is the second of two devoted to proving Theorem 1.1. Here we show that for measures supported away from certain divisors, a quantitative version of Theorem 1.1 holds.
We will now assume that $k_{0}\in\mathbb{N}$ is big enough so that $n_{k_{0}}\geq 2$. We fix two linearly independent sections $s_{1},s_{2}\in H^{0}(X,L^{\otimes k_{0}})$, and consider the subvariety $Y\mathrel{\mathop{\ordinarycolon}}={\textrm{div}}(s_{2})$. We fix a Borel measure $\mu_{0}$ on $X$ whose support is disjoint from $Y$. If the support of $\mu_{0}$ is not an analytic subset, then the form defined as in (1.1), but for $\mu_{0}$ instead of $\mu$, is positive definite, and hence the associated Bergman kernel, $B_{k,0}(x,y)\in L^{k}_{x}\otimes(L^{k}_{y})^{*}$, can be defined. In general, one can define $B_{k,0}(x,y)$ by considering an orthonormal basis of a vector subspace $E_{k}\subset H^{0}(X,L^{\otimes k})$ of maximal dimension, so that the restriction map $E_{k}\to L^{2}(\mu)$ is injective. While $|B_{k,0}(x,y)|^{2}_{(h^{L})^{k}}$ might depend on the choice of $E_{k}$, an easy verification shows that the resulting measure $|B_{k,0}(x,y)|^{2}_{(h^{L})^{k}}d\mu_{0}(x)d\mu_{0}(y)$ is independent of it. The principal result of this section — obtained without requiring any condition on $\mu_{0}$ other than that its support be disjoint from $Y$ — is the following.
Theorem 2.8.
There is $C>0$, so that for any $k\in\mathbb{N}^{*}$, we have
Remark 2.9.
b) In the setting of orthogonal polynomials described in (1.4), one can consider $s_{2}\mathrel{\mathop{\ordinarycolon}}=\sigma^{2}$ and $s_{1}=z\cdot\sigma^{2}$, where $\sigma\in H^{0}(\mathbb{P}^{1},\mathscr{O}(1))$ is as described after (1.4). Then for any compact $K\subset\mathbb{C}\times\mathbb{C}\subset\mathbb{P}^{1}\times\mathbb{P}^{1}$, not intersecting the diagonal, the above $C_{1}$ exists. We conclude that for arbitrary compactly supported measures $\mu$ on $\mathbb{C}$ with non-finite support, the estimate (1.7) holds. We note that when the support of $\mu$ lies inside of the real axis, one can give a shorter proof of (1.7) based on the Christoffel-Darboux formula, see in particular [19, (2.13)].
Proof.
Our proof is inspired by the the proof of [8, Theorem 17] by Berman-Ortega-Cerdá in the sense that the estimate (2.36) is established by studying the Hilbert-Schmidt norm of a certain sequence of operators. However, in contrast to [8], our operators are not of Toeplitz type. This allows us to bypass – at the expense of a more delicate analysis of the associated Schwartz kernel – the assumption present in [8] that the support of the measure is totally real. A somewhat similar sequence of operators was used in the context of orthogonal polynomials by Simon [58].
Let us verify that the operators $S_{k}$, $k\in\mathbb{N}$, are uniformly bounded, i.e. there is $C>0$, so that $\|S_{k}\|\leq C$, where $\|\cdot\|$ is the operator norm. Since the measure $\mu_{0}$ is supported away from $Y$, the function $f$ is uniformly bounded on its support. The operator $S_{k}$ is given by a composition of a Bergman projection (which is of unit norm) and the commutator of the Bergman projection and the multiplication operator by $f$ (which is of bounded norm, independent of $k\in\mathbb{N}$). Hence, there is $C>0$, so that $\|S_{k}\|\leq C$, for any $k\in\mathbb{N}$.
As a conclusion, we deduce the following bound
$\|S_{k}\|_{{\textrm{HS}}}^{2}\leq C\cdot(n_{k}-n_{k-k_{0}}),$ (2.38)
We expand the integrand, apply the reproducing property of the Bergman kernel to eliminate two of the four kernel factors in each of the resulting terms as follows
$\iint B_{k,0}(x,z)B_{k,0}(z,y)B_{k,0}(w,x)B_{k,0}(y,w)\cdot f(x)\overline{f}(x)\cdot d\mu_{0}(z)d\mu_{0}(w)$ (2.42)
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad=\big|B_{k,0}(x,y)\big|^{2}_{(h^{L})^{k}}\cdot f(x)\overline{f}(x),$
$\iint B_{k,0}(x,z)B_{k,0}(z,y)B_{k,0}(w,x)B_{k,0}(y,w)\cdot f(x)\overline{f}(w)\cdot d\mu_{0}(z)d\mu_{0}(y)$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad=\big|B_{k,0}(x,w)\big|^{2}_{(h^{L})^{k}}\cdot f(x)\overline{f}(w),$
$\iint B_{k,0}(x,z)B_{k,0}(z,y)B_{k,0}(w,x)B_{k,0}(y,w)\cdot f(z)\overline{f}(x)\cdot d\mu_{0}(w)d\mu_{0}(y)$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad=\big|B_{k,0}(x,z)\big|^{2}_{(h^{L})^{k}}\cdot f(z)\overline{f}(x),$
$\iint B_{k,0}(x,z)B_{k,0}(z,y)B_{k,0}(w,x)B_{k,0}(y,w)\cdot f(z)\overline{f}(w)\cdot d\mu_{0}(x)d\mu_{0}(y)$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad=\big|B_{k,0}(z,w)\big|^{2}_{(h^{L})^{k}}\cdot f(z)\overline{f}(w).$
By renaming the variables, we obtain
$\iint\big|S_{k}(x,y)\big|^{2}_{(h^{L})^{k}}\cdot d\mu_{0}(x)d\mu_{0}(y)\\ =\frac{1}{2}\iint\big|B_{k,0}(x,y)\big|^{2}_{(h^{L})^{k}}\cdot\big|f(x)-f(y)\big|^{2}\cdot d\mu_{0}(x)d\mu_{0}(y),$ (2.43)
which finishes the proof of Theorem 2.8 by (2.39). ∎
We are now finally in a position to prove the main result of this paper.
Proof of Theorem 1.1.
[…]
We now consider a subset $Y\mathrel{\mathop{\ordinarycolon}}=\Sigma\cup{\textrm{div}}(s_{1})\cup\cdots\cup{\textrm{div}}(s_{n_{k}})$, where $s_{i}$, $i=1,\ldots,n_{k}$, is a fixed basis of $H^{0}(X,L^{\otimes k})$. The set $Y$ is a subvariety. Hence, according to Theorem 2.3 and Proposition 2.4, it suffices to establish Theorem 1.1 for measures $\mu$ which are supported away from $Y$. Note, however, that for such measures $\mu$, for a fixed $K\subset X\times X$ not intersecting the diagonal, there is $C>0$, so that for $K^{\prime}\mathrel{\mathop{\ordinarycolon}}=K\cap{\rm{supp}}(\mu)\times{\rm{supp}}(\mu)$, for any $x,y\in K^{\prime}$, we have
3 On the weak limit of the diagonal Bergman measures
This section is mostly devoted to the following problem: which measures arise as subsequential limits of diagonal Bergman measures? As a consequence of our study, we will derive Theorem 2.3 in Section 3.1, modulo a statement that is established in Section 3.2. In Theorem 3.3, we then present a generalization of Theorem 2.3 and discuss a related open problem. Finally, Section 3.4 is devoted to the proof of Theorem 1.5. We retain the notation from Introduction and assume throughout the rest of the section that the measure $\mu$ does not give full mass to pluripolar subsets.
3.1 Diagonal Bergman measures put no mass on subvarieties
The primary objective of this section is to demonstrate that every subsequential limit of the diagonal Bergman measures vanishes on subvarieties, i.e. to show Theorem 2.3. Remark that when $\mu$ is Bernstein-Markov with non-pluripolar support, Theorem 2.3 follows from Theorem 2.2, as equilibrium measure puts no mass on pluripolar subsets (in particular on analytic subsets). In order to prove Theorem 2.3 in full generality, we consider the so-called partial Bergman kernels associated with $Y$, denoted by $B_{k}^{\epsilon,Y}(x,y)\in L^{k}_{x}\otimes(L^{k}_{y})^{*}$ for $\epsilon>0$, and defined as
@@ 3.1 Diagonal Bergman measures put no mass on subvarieties
where $s_{i}$, $i=1,\ldots,n_{k}(\epsilon)$ is an orthonormal basis of $(H^{0}(X,L^{\otimes k}\otimes\mathcal{J}_{Y}^{\lceil\epsilon k\rceil}),{\textrm{Hilb}}_{k}(h^{L},\mu))$, $n_{k}(\epsilon)\mathrel{\mathop{\ordinarycolon}}=\dim H^{0}(X,L^{\otimes k}\otimes\mathcal{J}_{Y}^{\lceil\epsilon k\rceil})$, and $\mathcal{J}_{Y}$ is the sheaf of holomorphic sections vanishing along $Y$. The following result, out of independent interest, will be established in Section 3.2.
@@
For any $\epsilon,\delta>0$, there is an open neighborhood $U$ of $Y$, so that for any $x\in U$, $k\in\mathbb{N}$, we have $B_{k}^{\epsilon,Y}(x,x)\leq\exp(-\delta k)$.
@@
When $\mu$ is a volume form, Theorem 3.1 follows from more precise estimates of Berman [10, Lemma 4.1].
@@
By Fujita’s theorem [35], the above upper limit is actually a limit. The definition of the volume functional naturally extends to the $\mathbb{Q}$-line bundles, and the resulting functional extends continuously to the full Néron-Severi space $NS^{1}(X)$, cf. [40, Theorem 2.2.44].
@@
By choosing $\epsilon\mathrel{\mathop{\ordinarycolon}}=\epsilon^{\prime}$ and $\delta\mathrel{\mathop{\ordinarycolon}}=1$ in Theorem 3.1, we get a neighborhood $U$ of $Y$, so that for any $x\in U$, $k\in\mathbb{N}$, we have
$B_{k}^{\epsilon^{\prime},Y}(x,x)\leq\exp(-k).$ (3.5)
By writing $B_{k}(x,x)=\big(B_{k}(x,x)-B_{k}^{\epsilon^{\prime},Y}(x,x)\big)+B_{k}^{\epsilon^{\prime},Y}(x,x)$, we obtain
$\int_{U}B_{k}(x,x)\cdot d\mu(x)\leq\int_{U}\big(B_{k}(x,x)-B_{k}^{\epsilon^{\prime},Y}(x,x)\big)\cdot d\mu(x)+\exp(-k).$ (3.6)
Due to the positivity of the integrand, we have
$\int_{U}\big(B_{k}(x,x)-B_{k}^{\epsilon^{\prime},Y}(x,x)\big)\cdot d\mu(x)\leq\int_{X}\big(B_{k}(x,x)-B_{k}^{\epsilon^{\prime},Y}(x,x)\big)\cdot d\mu(x)$ (3.7)
Using the elementary identities
@@
together with (3.4), we however deduce
$\int_{X}\big(B_{k}(x,x)-B_{k}^{\epsilon^{\prime},Y}(x,x)\big)\cdot d\mu(x)\leq\epsilon\cdot n_{k}.$ (3.9)
A combination of (3.6) and (3.9) yields (2.20). ∎
3.2 Forbidden region for the partial Bergman kernels
The main goal of this section is to establish that the partial Bergman kernels decrease exponentially in the neighborhood of the subvariety, that is, to prove Theorem 3.1.
To simplify the presentation, we claim that it suffices to establish Theorem 3.1 under the assumption that $\mathcal{J}_{Y}=\mathscr{O}_{\hat{X}}(-E)$ for some effective divisor $E$ on $X$. To see this, let $\pi\mathrel{\mathop{\ordinarycolon}}\hat{X}\to X$ denote the blow-up of $X$ along $\mathcal{J}_{Y}$, and let $\mathscr{O}_{\hat{X}}(-E)\mathrel{\mathop{\ordinarycolon}}=\pi^{*}\mathcal{J}_{Y}$. If $\hat{X}$ is not smooth, we perform further blow-ups to resolve its singularities. By an abuse of notation, we denote the resolution by $\hat{X}$. Let $\hat{\mu}$ be a measure on $\hat{X}$ such that $\pi_{*}\hat{\mu}=\mu$. Since $\mu$ does not give full mass to pluripolar subsets and $\pi$ is a biholomorphism outside an analytic (hence pluripolar) subset, the measure $\hat{\mu}$ does not give full mass to pluripolar subsets. The isomorphism (3.3) is then an isometry with respect to the induced $L^{2}$-norms. In particular, the associated diagonal Bergman measures $\hat{\mu}_{k}^{\mathrm{Berg},\Delta}$ on $\hat{X}$ and $\mu_{k}^{\mathrm{Berg},\Delta}$ on $X$ satisfy the compatibility relation $\pi_{*}\hat{\mu}_{k}^{\mathrm{Berg},\Delta}=\mu_{k}^{\mathrm{Berg},\Delta}$. It follows that it is enough to prove Theorem 3.1 on $\hat{X}$. This allows us to assume from now on that $\mathcal{J}_{Y}=\mathscr{O}_{\hat{X}}(-E)$ for some effective divisor $E$ on $X$.
Our proof of Theorem 3.1 will be based on the following three auxiliary statements. For the first one, we use the following notation: for a subset $K\subset X$, we denote by ${\textrm{Ban}}_{k}^{\infty}(K,h^{L})$ the $\sup$-seminorm on $H^{0}(X,L^{\otimes k})$ induced by $h^{L}$ and evaluated over $K$. For brevity, we denote ${\textrm{Ban}}_{k}^{\infty}(X,h^{L})$ by ${\textrm{Ban}}_{k}^{\infty}(h^{L})$.
Proposition 3.3.